### v2

# ImFace++: A Sophisticated Nonlinear 3D Morphable Face Model with Implicit Neural Representations

Abstract
Accurate representations of 3D faces are of paramount importance in various computer vision and graphics applications. However, the challenges persist due to the limitations imposed by data discretization and model linearity, which hinder the precise capture of identity and expression clues in current studies. This paper presents a novel 3D morphable face model, named ImFace++, to learn a sophisticated and continuous space with implicit neural representations. ImFace++ first constructs two explicitly disentangled deformation fields to model complex shapes associated with identities and expressions, respectively, which simultaneously facilitate the automatic learning of correspondences across diverse facial shapes. To capture more sophisticated facial details, a refinement displacement field within the template space is further incorporated, enabling a fine-grained learning of individual-specific facial details. Furthermore, a Neural Blend-Field is designed to reinforce the representation capabilities through adaptive blending of an array of local fields. In addition to ImFace++, we have devised an improved learning strategy to extend expression embeddings, allowing for a broader range of expression variations. Comprehensive qualitative and quantitative evaluations demonstrate that ImFace++ significantly advances the state-of-the-art in terms of both face reconstruction fidelity and correspondence accuracy.
Index Terms:
3D Morphable Face Models, Face Geometry, Implicit Neural Representations, Implicit Displacement Fields.
1 Introduction
3D Morphable Face Models (3DMMs) are well-reputed statistical models, established by learning techniques upon prior distributions of facial shapes and textures from a set of samples with dense correspondence, aiming at rendering realistic faces of a high variety. Since a morphable representation is unique across different downstream tasks where the geometry and appearance are separately controllable, 3DMMs are pervasively exploited in many face analysis applications in the field of computer vision, computer graphics, biometrics, and medical imaging [4, 2, 3, 5].
In 3DMMs, the most fundamental issue lies in the way to generate latent morphable representations, and during the past two decades, along with data improvement in scale, diversity, and quality [8, 11, 6, 41], remarkable progresses have been achieved. The methods are initially linear model based [1, 57, 58] and further extended to multilinear model based [9, 7, 56], where different modes are individually encoded. Unfortunately, for the relatively limited representation ability of linear models, these methods are not so competent at handling the cases with complicated variations, e.g. , exaggerated expressions. In the era of deep learning, a number of nonlinear models have been investigated with the input of 2D images [12, 19] or 3D meshes [14, 15, 16, 17, 44] by using Convolutional Neural Networks or Graph Neural Networks. They indeed deliver performance gains; however, restricted by the resolution of discrete representing strategies on input data, facial priors are not sufficiently captured, incurring loss of shape details. Besides, all current methods are dependent on the preposed procedure of point-to-point correspondence [10, 59, 60, 43], but face registration itself remains challenging.
Figure 1: ImFace++ effectively encodes the intricate face variations associated with identities and expressions by two explicitly disentangled deformation fields with respect to a template face. Leveraging the established correspondence, it applies an additional refinement displacement field to capture high-frequency details in the individual-independent template space. This approach yields sophisticated and implicit representations that are highly adaptable for 3D face modeling.
Recently, several studies have illuminated the potential of Implicit Neural Representations (INRs) [20, 21, 22, 26, 27] in the precise modeling of 3D geometries. INRs describe input observations as low-dimensional shape embeddings and estimate the Signed Distance Function (SDF) or the occupancy values of query points, allowing for the definition of surfaces with arbitrary resolution and topology through isocontours. Such continuous representations offer advantages over the traditional discrete ones like voxels, point clouds, and meshes, and have reported substantial success in shape reconstruction [62, 61, 28, 63], as well as surface registration [23, 25, 24]. Such an advantage suggests an alternative to 3DMM that can fulfill accurate correspondence and geometry deformation modeling in a unified network. Despite that, applying INRs to face modeling presents unique challenges. Unlike the objects with distinct shapes and limited non-rigid variations, such as indoor scenes and human bodies, facial surfaces are inherently similar but also encompass intricate deformations arising from the interweaving of multiple identities and rich expressions. Current INR methods struggle to precisely capture and disentangle these intricacies, as evidenced by the preliminary attempt  [33].
On the other side, it is essential to emphasize the faithful reproduction of facial details. In the context of geometry modeling, there has been a growing interest in hybrid approaches that combine explicit and implicit representations to harness the strengths of both techniques [74, 72, 69], aiming to address the spectral bias [65] issue often observed in neural networks and learn more sophisticated geometry clues. However, incorporating explicit 3D representations into INRs can introduce discontinuities, potentially leading to training instability and over-fitting [91], especially when applied to complex structures like facial surfaces. An alternative research trend focuses on incorporating computer graphics techniques, such as displacement mapping [82, 83], into INRs to encode rich details. While these methods have shown remarkable performance in reconstructing single static scenes, the challenge remains in developing INR models capable of effectively capturing dynamically changing high-frequency details across multiple facial shapes, such as wrinkles and dimples.
In response to the aforementioned challenges, this paper presents a novel two-stage 3D face morphable model, namely ImFace++, which substantially upgrades conventional 3DMMs by integrating customized INRs. In its initial stage, ImFace++ tackles nonlinear facial deformation modeling and establishment of correspondence among different individuals by creating separate INR sub-networks, which explicitly disentangle shape morphs into two deformation fields for identity and expression, respectively (as Fig. 1 shows). The template shape is defined by a deep SDF, therefore eliminating the need for the face registration phase required by numerical solutions. To build more sophisticated 3DMMs, ImFace++ further incorporates a refinement displacement field (RDF) in the second stage to capture high-frequency geometry details present in various faces. Notably, this stage for detail modeling is specifically tailored to adapt to the dynamic nature of facial expressions. By leveraging the established correspondences, RDF is learned within the individual-independent template space characterized by its relative stability, allowing the model to concentrate on encoding subtle details. Overall, by learning the disentangled deformation fields and RDF, ImFace++ accurately represents inter-individual changes and fine-grained clues while retaining the flexibility for its application in related tasks.
Inspired by linear blend skinning [37], we have further introduced a Neural Blend-Field to decompose the entire facial deformation or geometry into semantically meaningful regions encoded by an array of local implicit functions and adaptively blend them through a lightweight module, leading to more advanced representations with reduced parameter overhead. In addition to ImFace++, an improved embedding learning strategy is devised to extend the latent space of expressions to allow a more diverse expression range.
More concisely, this study makes the following contributions:
  • A novel INR-based 3DMM, namely ImFace++, which encodes complex facial shape variations associated with identity and expression through explicitly disentangled deformation fields. It enables the learning of flexible and semantically meaningful representations.
  • A refinement displacement field (RDF) within the individual-independent template space. Leveraging the established correspondence, this field faithfully encodes high-frequency details, thereby enhancing the representation capabilities of ImFace++.
  • A Neural Blend-Field, which enables fine-grained learning by adaptively blending an array of local implicit functions.
  • Extensive experiments evidence that ImFace++ excels in synthesizing high-quality 3D faces with intricate details, outperforming the state-of-the-art alternatives.
A preliminary version of this study was ImFace, previously published in [87]. This paper significantly improves it in the following aspects: (i) We extend the model to a two-stage framework. To sufficiently capture facial priors, the upgraded model effectively leverages the correspondences between different individuals, achieved by employing disentangled deformation fields in the first stage as in ImFace [87]; it further investigates a newly designed refinement field to progressively model face geometry, significantly alleviating the loss of shape details while relieving the burden of model learning. (ii) We update the model to reconstruct and synthesize 3D faces spanning a wide range of facial expressions and it renders more vivid results. We have also refined evaluations in both the observed and latent representation spaces, along with more comprehensive discussions. (iii) To further validate the proposed approach, we conduct comparisons with other INR-based methods [33, 66] in terms of 3D face reconstruction and face correspondence. (iv) We showcase applications in various scenarios, including facial expression editing and detail transfer, to demonstrate the adaptability of ImFace++ in the realm of 3D face modeling.
The rest of this paper is organized as follows. Section 2 reviews related work on 3DMMs and INRs. Section 3 elaborates on the proposed ImFace++ method. Section 4 presents and analyzes the experimental results using the FaceScape database, followed by Section 5, which concludes this paper and provides insights for future perspectives.
2 Related Work
The published studies closely related to our study can be primarily summarized into: (i) 3D Morphable Models, (ii) Implicit Neural Representations, offering innovative methods for capturing complex 3D structures, and (iii) Fine-Grained Neural Representations for Geometry Modeling, delving into studies that specifically address the creation and application of fine-grained neural representations for geometry modeling. In the following, we provide a concise review of the relevant literature.
2.1 3D Morphable Face Models
The concept of 3D Morphable Face Models (3DMMs) was first introduced by Blanz and Vetter [1] as a foundational framework for facial representation. This approach involved registering a known template mesh to all training scans using the Non-rigid Iterative Closest Point (NICP) algorithm [40]. Principal Components Analysis (PCA) was subsequently employed to create prior face distributions. To account for identity-dependent expressions, 3DMMs were extended to multilinear models [9, 7, 56]. Over time, significant progress has been made aided by improvements in data quality [8, 11, 6, 41]. Notably, the Facial LAndmarks and More Expressions (FLAME) model [6] was introduced as an expressive model, capable of controlling facial expressions through a combination of jaw articulation and linear expression blendshapes. Trained on a comprehensive 3D facial dataset, including D3DFACS [42], FLAME delivered remarkable results, although it faced challenges in accurately capturing nonlinear facial deformations.
The development of deep neural networks has significantly accelerated the development of more potent nonlinear 3DMMs. A number of models have been trained directly from 2D images [12, 19, 18], but they often struggled to capture high-fidelity and fine-grained details. This limitation is attributed to the inherent challenges posed by the low resolution of input images in solving this ill-posed inverse problem. Recognizing the need to better leverage 3D face scans, Bagautdinov and colleagues [14] devised a method for mapping 3D meshes to the 2D space. Subsequently, several other studies [44, 15, 17, 16] directly pursued the learning of 3DMMs from meshes, employing spectral or spiral convolutions. While promising, these neural networks were established on discrete 3D representations, which posed limitations when confronted with intricate facial deformations.
For a more comprehensive survey on 3DMMs, readers are encouraged to refer to [13].
2.2 Implicit Neural Representations
In recent years, implicit neural functions have emerged as powerful and effective representations for 3D geometry [20, 21, 22, 26, 27, 45]. These approaches have revolutionized the field by enabling continuous and highly expressive representations of geometric structures without the need for discretization. To address the complexities of shape variations and correspondence relationships, some methods have further incorporated an additional implicit deformation latent space [23, 24].
A number of studies have focused on applying INRs to model human heads or bodies [35, 34, 33, 32, 31, 36, 30, 47, 66, 88, 89, 90, 92, 93, 95, 94]. In particular, NPMs [95] and GNPM [94] have introduced approaches employing two neural networks to separately model body shape and pose spaces. Models such as H3D-Net [34] and SIRA++[92] have learned implicit priors for 3D head reconstruction. i3DMM[33] made significant strides by pioneering a generative implicit 3D morphable model tailored for human heads. However, utilizing neural networks to generate fine-grained details from raw signals has often resulted in over-smoothed or implausible facial shapes in these studies. A concurrent work, NPHM [66], has devised a parametric head model using local implicit fields, similar to our approach. It employed local INRs for identity representation and a globally conditioned INR for modeling expression deformations. Nonetheless, its ability for capturing high-frequency facial details was constrained by the globally conditioned neural network architecture it employed. Another series of works [88, 89, 90, 93] have concentrated on modeling facial appearance, leveraging techniques like differential volume rendering [49] and auto-decoding [20] to learn morphable radiance fields. However, these models represented the underlying geometry using density, leading to noisy geometry approximations.
Figure 2: ImFace++ overview. The model is constructed through two-stage learning. (a) The first stage primarily focuses on nonlinear facial deformation modeling and correspondence establishment between different faces. The network comprises three Mini-Nets blocks designed to explicitly disentangle shape morphs into separate deformation fields, where the Expression and Identity Mini-Nets blocks are associated with expression and identity deformations, respectively. Meanwhile, the Template Mini-Nets block learns the SDF of a template face shape. (b) With the correspondence established, ImFace++ proceeds to employ a Detail Mini-Nets block in the second stage, responsible for learning refinement displacement field that encodes high-frequency geometry details.
2.3 Fine-Grained Neural Representations for Geometry Modeling
Recent research [65, 67, 68] has brought attention to a persistent challenge in neural networks that the learned models tend to favor the retention of low-frequency signals while neglecting high-frequency information, known as the spectral bias [65]. In the context of geometry modeling, recent advancements have introduced hybrid approaches that combine implicit and explicit 3D representations, e.g. , voxels [69, 70, 71], tri-plane [72, 73] and hash-grids [74, 75, 76], to mitigate the issue above. In particular, Voxurf [71] employed voxel-based representations combined with an implicit function defined by a shallow MLP to represent a scene. It also introduced hierarchical geometry features to enhance perception across voxels. EG3D [72] presented a memory-efficient tri-plane architecture that scales efficiently with resolution, yielding higher capacity for handling complex 3D data. TensoRF [77] adopted a unique approach to decompose the 3D volume into vector-matrix outer products, leading to improved rendering quality. Additionally, INGP [74] combined voxel grids and multi-resolution hash tables to create compact and adaptive data representations. This innovative approach resulted in significant enhancements in both rendering quality and processing speed. However, it is worth noting that while these hybrid 3D representations offer advantages in various applications, using them as 3DMMs can be challenging due to issues related to the discontinuity of the underlying explicit 3D representations. Such discontinuity may result in instability and incorrect training outcomes, particularly when modeling complex surfaces such as facial shapes.
Another emerging research trend seeks to bolster the intrinsic capabilities of neural networks for improved details. Inspired by the natural language processing field, NeRF [49] utilized position encoding to render finer-grained details. In addition, SIREN [78] and GARF [79] introduced innovative network activation functions, utilizing Sin and Gaussian functions, respectively. However, training these networks can be a challenging endeavor due to instability and susceptibility to over-fitting. To overcome this, FreeNeRF [80] and Nerfies [81] proposed to use a coarse-to-fine learning strategy. Furthermore, IDF [82] and HF-NeuS [83] extended the conventional explicit and discrete displacement mapping into ${\mathbb{R}}^{3}$ domain to capture high-frequency details within single static scenes. Nevertheless, there is relatively limited exploration into the modeling of dynamic high-frequency facial details across different individuals. To enhance the ability of INRs to build 3DMMs, our method employs a decomposition of the learning process. This approach begins by automatically learning deformations linked to identity and expression while simultaneously establishing correspondences between different individuals. It then learns to encode dynamic high-frequency details within the established template space, significantly facilitating the representation capability.
3 Method
We take advantages of INRs to learn a sophisticated nonlinear 3DMM. The fundamental idea of INRs is to train a neural network to fit a continuous function $f$, which implicitly represents surfaces through level-sets. The function can be defined in various formats, e.g. , occupancy [21], SDF [20], or UDF [38]. We exploit a deep SDF conditioned on the latent embeddings associated with various facial characteristics for comprehensive facial representations. It outputs the signed distance $s$ from a query point:
$f:({\mathbf{p}},{\mathbf{z}}_{exp},{\mathbf{z}}_{id},{\mathbf{z}}_{detail})\in{\mathbb{R}}^{3+d_{exp}+d_{id}+d_{detail}}\mapsto s\in{\mathbb{R}},$ (1)
where ${\mathbf{p}}\in{\mathbb{R}}^{3}$ is the coordinate of the query point in the 3D space, ${\mathbf{z}}_{exp}$, ${\mathbf{z}}_{id}$, and ${\mathbf{z}}_{detail}$ denote the expression, identity and high-frequency detail embeddings of a given face, respectively.
Our goal is to learn a neural network to parameterize $f$, making it satisfy the genuine facial shape priors. As shown in Fig. 2, the proposed network for ImFace++ is constructed by a series of Mini-Nets blocks that are progressively learned. In the first stage, ImFace++ employs the Expression and Identity Mini-Nets blocks to learn separate deformation fields associated with expression- and identity-variation, respectively. Meanwhile, a Template Mini-Nets block is used to learn a signed distance field of a template face shape. Such explicitly disentangled architecture ensures that the intricate expression deformations and inter-individual differences can be accurately modeled (Sec. 3.1). The second stage complements the former by incorporating a Detail Mini-Nets block to learn a refinement displacement field (RDF) within the template space, which faithfully encodes facial geometry details (see Sec. 3.2). The network parameters and facial embeddings are jointly learned. After convergence, a face scan be disentangled by deforming it from its observed state to the template.
All the aforementioned fields are implemented by a shared Mini-Nets architecture, where the entire facial deformation or geometry is decomposed into an array of semantically meaningful parts and encoded by a set of local field functions so that intricate details can be sufficiently captured. A lightweight module conditioned on the query point position, i.e, Fusion Network, is stacked at the end of the Mini-Nets block to adaptively blend the local fields. As such, an elaborate Neural Blend-Field is achieved (see Sec. 3.3). Additionally, an improved embedding learning strategy is designed, allowing more diverse details of expression (refer to Sec. 3.4). Detailed explanations of the model designs are provided as follows.
3.1 Disentangled Deformation Fields
The proposed method initially focuses on non-linear facial deformation modeling and the establishment of correspondences among different faces. This is accomplished through three Mini-Nets blocks as previously outlined. As Fig. 2 (a) shows, we collectively refer to this process as Disentangled Deformation Fields modeling and the learned SDF can reproduce a base observation space. The core components serve distinct purposes, with slight structural adaptations as follows:
3.1.1 Expression Mini-Nets (ExpNet)
The facial deformations incurred by expressions are represented by ExpNet $\mathcal{E}$, which learns an observation-to-canonical warping for every face scan:
$\mathcal{E}:({\mathbf{p}}_{b},{\mathbf{z}}_{exp},l)\mapsto{\mathbf{p}}^{\prime}\in{\mathbb{R}}^{3},$ (2)
where ${\mathbf{p}}_{b}$ denotes a query point in the base observation space, and $l\in{\mathbb{R}}^{k\times 3}$ denotes $k$ 3D landmarks on the observed face generated by a Landmark-Net $\eta:({\mathbf{z}}_{exp},{\mathbf{z}}_{id})\mapsto l$, which is introduced to localize the query point ${\mathbf{p}}_{b}$ in the Neural Blend-Field. This point is deformed by $\mathcal{E}$ to ${\mathbf{p}}^{\prime}$ in the person-specific canonical space, which represents faces with a neutral expression.
3.1.2 Identity Mini-Nets (IDNet)
To model shape morphs among individuals, IDNet $\mathcal{I}$ further warps the canonical space to a template shape space shared by all faces:
$\mathcal{I}:({\mathbf{p}}^{\prime},{\mathbf{z}}_{id},l^{\prime})\mapsto({\mathbf{p}}^{\prime\prime},\delta)\in{\mathbb{R}}^{3}\times{\mathbb{R}},$ (3)
where $l^{\prime}\in{\mathbb{R}}^{k\times 3}$ denotes $k$ landmarks on the canonical face generated by another Landmark-Net conditioned only on the identity embedding, namely $\eta^{\prime}:{\mathbf{z}}_{id}\mapsto l^{\prime}$, where ${\mathbf{p}}^{\prime\prime}$ is the deformed point in the template space. To cope with the possible non-existent correspondences produced during preprocessing, $\mathcal{I}$ additionally predicts a residual term $\delta\in{\mathbb{R}}$ to correct the SDF value $s_{0}$ predicted by the subsequent TempNet, which is similar to [23].
3.1.3 Template Mini-Nets (TempNet)
TempNet $\mathcal{T}$ learns a signed distance field of the shared template face:
$\mathcal{T}:({\mathbf{p}}^{\prime\prime},l^{\prime\prime})\mapsto s_{0}\in{\mathbb{R}},$ (4)
where $l^{\prime\prime}\in{\mathbb{R}}^{k\times 3}$ denotes $k$ landmarks on the template face, which is averaged on the whole training set, and $s_{0}$ is the uncorrected SDF value. The final SDF value of a query point is calculated via $s=s_{0}+\delta$, and the modeling process of disentangled deformation fields can be collectively formulated as:
$\hat{f}({\mathbf{p}}_{b},{\mathbf{z}}_{exp},{\mathbf{z}}_{id})=\mathcal{T}(\mathcal{I}(\mathcal{E}({\mathbf{p}}_{b})))+\mathcal{I}_{\delta}(\mathcal{E}({\mathbf{p}_{b}})),$ (5)
For brevity, the latent embeddings and predicted landmarks are omitted. By doing so, we can automatically learn correspondence across different individuals by mapping the query point ${\mathbf{p}}_{b}$ to the template point ${\mathbf{p}^{\prime\prime}}$.
Figure 3: Neural Blend-Field. (a) The Mini-Nets block is a shared architecture, which decomposes an entire facial feature into semantically meaningful parts and encodes them by a set of local field functions. It is tailed by a Fusion Network for more comprehensive representations. (b) The Landmark-Net is introduced to softly partition the entire facial surface. (c) The Fusion Network is a lightweight module conditioned on the query point position, which adaptively blends the local field functions, resulting in an elaborate Neural Blend-Field.
3.2 Refinement Displacement Field within Template Space
Facial modeling assumes a pivotal role in enhancing the overall visual fidelity and realism of facial representations. The intricate nuances, such as fine lines, wrinkles, and subtle expressions, collectively contribute to the creation of a truly lifelike and convincing depiction of a face. In order to construct an advanced 3DMM, we have introduced a specialized refinement module designed to master high-frequency facial details. Drawing inspiration from the concept of Implicit Displacement Fields (IDF) proposed in [82], where high-frequency details are generated by offsetting a smooth base surface along its normal directions based on the displacement distances predicted from IDF, we have developed a Detail Mini-Nets block (DetailNet), denoted as $\mathcal{D}$, to learn a refinement displacement field (RDF), as depicted in Figure 2 (b). To account for the dynamic nature of human faces, instead of representing facial geometry details using a conventional IDF, we encode them within the individual-independent template space that exhibits relative stability, so that subtle changes can be well captured. This strategy is reminiscent of the concept of storing displacement information in a topology-consistent UV texture map.
To achieve this, we leverage the correspondences established during the Disentangled Deformation Fields modeling phase and RDF is learned in the later stage. Formally, its modeling process can be expressed as:
$\mathcal{D}:({\mathbf{p}}^{\prime\prime},{\mathbf{z}}_{detail},l^{\prime\prime})\mapsto{\mathbf{d}}\in{\mathbb{R}},$ (6)
where ${\mathbf{d}}$ denotes the displacement distance along the normal of an observed point ${\mathbf{p}}$, which tends to be overlooked in the first learning stage. It is worth noting that ${\mathbf{p}}^{\prime\prime}$ is the input of this function, ensuring that the subtle geometry details are learned in the template space. Such learning phase decomposition effectively isolates subtle displacements from identity and expression-related deformations, and the facial details can be readily applied to other facial shapes that share the same template space. Taking the detail displacement into consideration, the coordinates ${\mathbf{p}}_{b}$ in the aforementioned coarse base observation space represented by $\hat{f}$ can be corrected as follows:
${\mathbf{p}}^{\prime}_{b}={\mathbf{p}}+\chi(\hat{f}({\mathbf{p}}))\mathcal{D}({\mathbf{p}}^{\prime\prime},{\mathbf{z}}_{detail},l^{\prime\prime})\frac{\nabla\hat{f}({\mathbf{p}})}{||\nabla\hat{f}({\mathbf{p}})||},$ (7)
where $\chi$ denotes the attenuation function [82] and $\frac{\nabla\hat{f}({\mathbf{p}})}{||\nabla\hat{f}({\mathbf{p}})||}$ denotes the normalized normal of ${\mathbf{p}}$. Finally, for a face scan, its key components (i.e., expression, and identity) can be accurately disentangled by performing the following warping process:
$f({\mathbf{p}},{\mathbf{z}}_{exp},{\mathbf{z}}_{id},{\mathbf{z}}_{detail})=\mathcal{T}(\mathcal{I}(\mathcal{E}({\mathbf{p}}^{\prime}_{b})))+\mathcal{I}_{\delta}(\mathcal{E}({\mathbf{p}}^{\prime}_{b})).$ (8)
The proposed ImFace++ learns face morphs by disentangled deformation fields and RDF in a fine-grained and meaningful manner, ensuring that more diverse and sophisticated facial deformations, correspondence, and details can be accurately learned.
3.3 Neural Blend-Field
The Mini-Nets block is a common architecture shared by the sub-networks $\mathcal{E}$, $\mathcal{I}$, $\mathcal{T}$, and $\mathcal{D}$. As shown in Fig. 3, it learns a continuous field function ${\bm{\psi}}:{\mathbf{x}}\in{\mathbb{R}}^{3}\mapsto v$ (e.g. , deformation, displacement distance or signed distance value), to produce a Neural Blend-Field for comprehensive face representations. In particular, to overcome the limited expressivity of a single network, we decompose a face into a set of semantically meaningful local regions and learn $v$ individually before blending. Such design is inspired by the recent INRs study [48] on human body, which introduces the linear blend skinning algorithm [37] to make the network learn from separate transformations of body parts. To better represent detailed facial shapes, we replace the constant transformation term in the original linear blend skinning algorithm with ${\bm{\psi}}_{n}({\mathbf{x}}-l_{n})$, and define the Neural Blend-Field as:
$v={\bm{\psi}}({\mathbf{x}})=\sum_{n=1}^{k}{{\it w}_{n}({\mathbf{x}}){\bm{\psi}}_{n}({\mathbf{x}}-l_{n})},\vspace{-2mm}$ (9)
where $l_{n}$ is a parameter that describes the $n$-th local region, ${\it w}_{n}({\mathbf{x}})$ is the $n$-th blend weight, ${\bm{\psi}}_{n}({\mathbf{x}}-l_{n})$ is the corresponding local field, and $k$ is the number of local regions. In this way, the blending is performed on a series of local fields, rather than calculating a weighted average of the output values $v$ of some fixed positions, leading to more powerful representation capability in handling complicated local features.
Specifically, five landmarks located at the outer eye corners, mouth corners, and nose tip are utilized to describe the local regions $(l_{n}\in{\mathbb{R}}^{3})_{n=1}^{5}$, which is predicted by Landmark-Net, as illustrated in Fig. 3 (b). As Fig. 3 (a) shows, each region is assigned a tiny MLP with sinusoidal activations [45] to generate the local field, denoted as ${\bm{\psi}}_{n}$. To capture high-frequency local variations, we leverage sinusoidal positional encoding $\gamma$ [49] on the coordinate ${\mathbf{x}}-l_{n}$. At the end of a Mini-Nets block, a lightweight Fusion Network conditioned on the absolute coordinate of input ${\mathbf{x}}$ is equipped, which is implemented by a 3-layer MLP with softmax to predict the blend weights $(w_{n}\in{\mathbb{R}}^{+})_{n=1}^{5}$, as illustrated in Fig. 3 (c).
3.3.1 Deformation Formulation
We formulate the deformation with a SE(3) field $({\bm{\omega}},{\mathbf{v}})\in{\mathbb{R}}^{6}$, where ${\bm{\omega}}\in so(3)$ is a rotate vector representing the screw axis and the angle of rotation. The deformed coordinates ${\mathbf{x}}^{\prime}$ can be calculated by ${e^{\bm{\omega}}}{\mathbf{x}}+{\mathbf{t}}$, where the rotation matrix ${e^{\bm{\omega}}}$ (exponential map form of Rodrigues’ formula) is written as:
${e^{\bm{\omega}}}={\bf I}+\frac{\sin\left\|{\bm{\omega}}\right\|}{\left\|{\bm{\omega}}\right\|}{\bm{\omega}}^{\wedge}+\frac{1-\cos\left\|{\bm{\omega}}\right\|}{{\left\|{\bm{\omega}}\right\|}^{2}}({\bm{\omega}}^{\wedge})^{2},$ (10)
and the translation ${\mathbf{t}}$ is formulated as:
${\mathbf{t}}=\left[{\bf I}+\frac{1-\cos\left\|{\bm{\omega}}\right\|}{{\left\|{\bm{\omega}}\right\|}^{2}}{\bm{\omega}}^{\wedge}+\frac{\left\|{\bm{\omega}}\right\|-\sin\left\|{\bm{\omega}}\right\|}{{\left\|{\bm{\omega}}\right\|}^{3}}({\bm{\omega}}^{\wedge})^{2}\right]{\mathbf{v}},$ (11)
where ${\bm{\omega}}^{\wedge}$ denotes the skew-symmetric matrix of ${\bm{\omega}}$. We exploit SE(3) to describe facial shape morphs for its superior capability in handling mandibular rotations and higher robustness to pose perturbations than the common translation deformation ${\mathbf{x}}^{\prime}={\mathbf{x}}+{\mathbf{t}}$.
3.3.2 Hyper Nets
To obtain a more compact and expressive latent space, we have introduced a meta-learning approach[45]. A Hyper Net $\phi_{n}$ is implemented by an MLP and it predicts the instance-specific parameters for ExpNet $\mathcal{E}$, IDNet $\mathcal{I}$ and DetailNet $\mathcal{D}$. It takes a latent code ${\mathbf{z}}$ as input and generates the parameters for the neurons in a Mini-Net ${\bm{\psi}}_{n}$ so that the learned facial representations possess a higher variety.
(a) Signed Distance Field
(b) Unsigned Distance Field
(c) SDF on pseudo watertight face
Figure 4: (a) SDF is able to represent closed shapes. (b) UDF is capable of representing an open surface but the gradient is discontinuous at the boundary, making it hard to be fitted by neural networks. (c) The proposed method generates pseudo watertight faces and restricts implicit functions on them, enabling implicit neural networks to learn geometry representations on 3D faces.
3.4 Improved Expression Embedding Learning
The auto-decoder framework proposed by [20] has been widely adopted in INRs to jointly learn embeddings and network parameters. In the previous attempt [33], each expression type is encoded by one embedding for attribute disentangling. Unfortunately, such an embedding is merely able to represent the average shape morph of an expression type, making the learned latent space fail to capture more diverse deformation details across individuals. To avoid the dilemma above, we improve the learning strategy by treating each face scan as a unique expression and generating a specific embedding for it. In this way, the latent space is significantly extended, which enables $\mathcal{E}$ to represent more fine-grained deformations. On the other side, there exists a potential failure mode that the identity properties are tangled into the expression space again, and $\mathcal{I}$ collapses to an identity mapping. To tackle this challenge, we suppress $\mathcal{E}$ when the current training sample is a neutral face, written as:
$\mathcal{E}({\mathbf{p}_{ne}},{\mathbf{z}}_{exp},l)\equiv{\mathbf{p}}_{ne},$ (12)
where ${\mathbf{p}_{ne}}$ denotes a point from a neutral face. By applying such a learning strategy, $\mathcal{I}$ and $\mathcal{T}$ jointly learn shape representations on neutral faces, and $\mathcal{E}$ focuses only on expression deformations. Moreover, only neutral labels are required during training, bypassing the need for dense expression labels.
3.5 Loss Functions
ImFace++ is trained with a set of loss functions to learn plausible facial shape representations, dense correspondence, and high-frequency details.
Reconstruction Loss. The basic SDF structure loss is applied to learn implicit fields:
${\cal L}_{sdf}^{i}={\lambda_{1}}\sum_{{\mathbf{p}}\in\Omega_{i}}\left|f({\mathbf{p}})-{\bar{s}}\right|+{\lambda_{2}}\sum_{{\mathbf{p}}\in\Omega_{i}}(1-\left\langle{\nabla f({\mathbf{p}})},{\mathbf{\bar{n}}}\right\rangle),$ (13)
where ${\bar{s}}$ and ${\mathbf{\bar{n}}}$ denote the ground-truth SDF values and the field gradients, respectively. $i$ denotes the index of sampled faces, and $\lambda$ indicates trade-off parameter. ${\Omega_{i}}$ is the sampling space of the scan $i$.
Eikonal Loss. To obtain reasonable fields throughout the network, multiple Eikonal losses are used to enforce the L-2 norm of spatial gradients to be unit:
${\cal L}_{eik}^{i}\!=$ ${\lambda_{3}}\!\sum_{{\mathbf{p}}\in\Omega_{i}}\!\big{(}|\|\nabla f({\mathbf{p}})\|\!-\!1|\!+\!|\|\nabla\mathcal{T}(\mathcal{I}({\mathbf{p}^{\prime}}))\|\!-\!1|\big{)},$ (14)
where ${\cal L}_{eik}^{i}$ serves the purpose of enabling the network to satisfy the Eikonal constraint[26] in both the observation and canonical spaces concurrently. This constraint plays a significant role in maintaining a sensible correspondence throughout facial deformations at all stages of the network.
Embedding Loss. It regularizes the embedding with a zero-mean Gaussian prior:
${\cal L}_{emb}^{i}={\lambda_{4}}\left({\|{\mathbf{z}}_{exp}\|}^{2}+{\|{\mathbf{z}}_{id}\|}^{2}\right)+{\lambda_{5}}{\|{\mathbf{z}}_{detail}\|}^{2}.$ (15)
Landmark Generation Loss. The ${\mathit{l}}_{1}$-loss is used to learn the Landmark-Nets $\eta$, $\eta^{\prime}$:
${\cal L}_{lmk_{g}}^{i}\!=\!{\lambda_{6}}\sum_{n=1}^{k}\left(|l_{n}-{\bar{l_{n}^{i}}}|+|l^{\prime}_{n}-{\bar{l_{n}^{\prime}}}|\right),$ (16)
where ${\bar{l^{i}}}$ denotes the $k$ labeled landmarks on sample $i$, ${\bar{l^{\prime}}}$ denotes the landmarks on the corresponding neutral face.
Landmark Consistency Loss. We exploit this loss to guide the deformed landmarks to be located at the corresponding positions on the ground-truth neutral and template faces for better correspondence performance:
${\cal L}_{lmk_{c}}^{i}\!=\!{\lambda_{7}}\sum_{n=1}^{m}\left(|\mathcal{E}(l_{n})-{\bar{l^{\prime}}}_{n}|+|\mathcal{I}(\mathcal{E}(l_{n}))-{l_{n}^{\prime\prime}}|\right),$ (17)
where $m$ denotes the point number of each batch sampled.
Residual Constraint. As in [23], we apply a penalty to $\delta$ to prevent it from excessively acquiring template face information, which could potentially compromise the quality of the morphable model. This penalty is implemented by:
${\cal L}_{res}^{i}={\lambda_{8}}\sum_{{\mathbf{p}}\in\Omega_{i}}|\delta({\mathbf{p}})|.$ (18)
Improved Expression Embedding Constraint. As described in Sec. 3.4, to avoid a potential failure mode that the identity properties are tangled into the expression space, we explicitly disentangle ExpNet $\mathcal{E}$ and IDNet $\mathcal{I}$ by:
${\cal L}_{imp}^{i}={\lambda_{9}}\sum_{{\mathbf{p}}\in\Omega_{i}}\begin{cases}||\mathcal{E}({\mathbf{p}})-{\mathbf{p}}||^{2},&{if\;i\;is\;a\;neutral\;face}\\ 0.&else\\ \end{cases}$
In the initial training phase, our focus lies on training the disentangled deformation fields, as elaborated in Section 3.1. This phase involves learning the coarse base SDF and establishing correspondences between various facial shapes. The training loss is formulated as follows:
${\cal L}_{\hat{f}}=\sum_{i}({\cal L}_{sdf}^{i}+{\cal L}_{eik}^{i}+{\cal L}_{emb}^{i}+{\cal L}_{lmk_{g}}^{i}+{\cal L}_{lmk_{c}}^{i}+{\cal L}_{res}^{i}+{\cal L}_{imp}^{i}).$ (19)
Subsequently, we freeze the sub-networks $\mathcal{E}$, $\mathcal{I}$, and $\mathcal{T}$ and proceed to train $\mathcal{D}$ to learn the refinement displacement field, by substituting $\hat{f}$ with $f$ in the following training loss:
${\cal L}_{f}=\sum_{i}({\cal L}_{sdf}^{i}+{\cal L}_{eik}^{i}+{\cal L}_{emb}^{i}).$ (20)
In alignment with the progressive training strategy in [82], we combine ${\cal L}_{f}$ and ${\cal L}_{\hat{f}}$ via $\kappa{\cal L}_{f}+(1-\kappa){\cal L}_{\hat{f}}$ using a weighted sum $\kappa=\frac{1}{2}(1+\cos(\pi\frac{t-T_{m}}{1-T_{m}}))$, where $T_{m}\in[0,1]$ and $t\in[T_{m},1]$ indicate the training percentile and the current training progress, respectively. We symmetrically adjust the learning rates, decreasing or increasing them in proportion to the loss weight.
At the testing phase, for each 3D face indexed by $j$, we minimize the following objective to obtain its latent embeddings and the reconstructed 3D face:
$\mathop{\arg\min}\limits_{{\mathbf{z}}_{exp},{\mathbf{z}}_{id},{\mathbf{z}}_{detail}}\sum_{j}({\cal L}_{sdf}^{j}+{\cal L}_{eik}^{j}+{\cal L}_{emb}^{j}).$ (21)
4 Experiments
We conduct thorough subjective and objective evaluations of ImFace++, along with ablation studies to validate the effectiveness of the specifically designed modules. Finally, we demonstrate various applications of ImFace++.
4.1 Dataset
FaceScape[41] is utilized for evaluation. It is a large-scale high quality 3D face dataset consisting of 938 individuals with 20 types of expressions. The data from 847 individuals are publicly accessible and we mainly use them for experiments. More specifically, 16,434 face scans from 831 individuals with nearly 20 expressions are sampled as the training set, and another set of 319 face scans from the remaining 16 individuals with nearly 20 expressions are used as our testing set. It is worth noting that, before conducting any data preprocessing, we make sure that the face scans include high-frequency details by making use of the displacement map provided by FaceScape.
Figure 5: Illustration of the preprocessing pipeline. The original face data from FaceScape [41] undergoes initial normalization and cropping. Hidden surfaces are subsequently removed. The Delaunay Triangulation Algorithm [50] is then employed for achieving watertight results. Lastly, we sample points and estimate the SDF and normals for training ImFace++.
Data Preprocessing. Given that neural networks are particularly adept at fitting functions that are differentiable across the entire domain, recent studies on implicit functions generally require watertight input data. While functions like UDF do not strictly necessitate watertightness, they exhibit non-differentiability when crossing a surface and may struggle to effectively handle fine details, as illustrated in Figure 4. To address this, we have developed an efficient preprocessing pipeline that transforms the input data into pseudo-watertight faces. This preprocessing step enables the establishment of a general SDF on these faces. By doing so, we ensure that the geometry and correspondence are learned with the same level of precision and fidelity as that of watertight objects.
As illustrated in Figure 5, the preprocessing consists of a series of steps: i) We rigidly align the facial scans to a frontal orientation using facial landmarks. Each mesh is then normalized to a standard unit of 10 cm. ii) The coordinate origin is defined at a point 4 cm behind the nose tip. A sphere with a radius of 10 cm is set as the sampling area, denoted as $S$. Mesh triangles outside of this sphere are cropped. iii) We apply the Ray-Triangle Intersection Algorithm [51] to remove hidden surfaces, such as the nasal and oral cavities. iv) Delaunay Triangulation[50] is performed on the x-y coordinates of the mesh to create an oriented and pseudo-watertight structure. v) Since FaceScape provides 3D face scans with uniform topology, we sample 13,008 points as landmarks within the defined sampling area $S$ of each scan to train ImFace++.
With the pseudo-watertight facial structures generated, we compute the SDF values through a distance transform. The sign of each sample is determined by the angle between its distance vector to the nearest surface and the positive z-axis direction. Values behind the facial surface are defined as negative. To ensure comprehensive coverage, we uniformly sample 250,000 points on each facial surface: 125,000 points near the facial surface with random displacements sampled from a normal distribution $\mathcal{N}(0,1cm)$, and 125,000 points within the sphere. For each of these points, we calculate their signed distances and gradient vectors. In the end, the sampled data is organized into triplets, denoted as ${({\bf p},{\bf\bar{n}},\bar{s})}$, representing the query point, gradient vectors, and signed distance value. These triplets serve as the input data for training ImFace++.
4.2 Implementation Details
Network Architecture. All Mini-Net blocks ${\bm{\psi}}_{n}$ used by ExpNet $\mathcal{E}$, IDNet $\mathcal{I}$, and TempNet $\mathcal{T}$, are implemented as Multi-Layer Perceptrons (MLPs) with 3 hidden layers, each consisting of 32-dimensional hidden features. In contrast, the Mini-Net block utilized by DetailNet $\mathcal{D}$ is implemented as MLPs with 3 hidden layers, comprising 64-dimensional hidden features. For the activation function, we employ sine activation in every Mini-Net $\psi_{n}$, and the parameters are initialized as described in [45]. The frequency hyperparameter $w$ in the sine activation function determines the capability of the neural network. We use a smaller value of $w=30$ for the disentangled deformation fields and a larger value of $w=60$ for DetailNet to enhance the capability of the networks. At the end of each Mini-Nets block, the Fusion Network is implemented as a 128-dimensional, 3-layer MLP with a softmax activation function. During training, the parameters of the $k$ Mini-Nets are generated by $k$ corresponding Hyper Nets. This approach is adopted to achieve a more expressive latent space, following the insights from [23] and [64]. The Hyper Nets, denoted as $\phi_{n}$, consist of 1-layer MLPs activated by the Rectified Linear Unit (ReLU) function, with the hidden layer dimensionality set to 32. The Landmark-Nets $\eta$ and $\eta^{\prime}$ are composed of three 256-dimensional fully connected layers. The identity, expression, and detail embeddings are each represented as 160-dimensional vectors.
Training Settings. The trade-off parameters $\lambda_{1},\cdots,\lambda_{9}$ to train the networks are set to $33e$ 3, $21e$ 2, $15e$ 1, $51e$ 5, $31e$ 3, $31e$ 3, $21e$ 2, $21e$ 2, and $41e$ 4, respectively. The model is trained with Adam in an end-to-end manner. We train the model for 1,500 epochs with a learning rate of 0.0001. The training phase takes around 4 days on 4 32G NVIDIA V100 GPUs with minibatches of size 48. We randomly sample $m=4096$ points from the total 13,008 points in each batch to train the landmark consistency loss. During testing, it takes about 8 hours to optimize 319 samples on a single GPU. We use Marching Cubes[86] to reconstruct facial surfaces from the signed distance field, where the voxel resolution is set to $256^{3}$. All the meshes are rendered by Pyrender[96].
4.3 Comparison with State-of-the-Art
4.3.1 Reconstruction
We employ the proposed ImFace++ model to fit face scans by optimizing Eq. (21) and then compare the reconstruction results with several existing methods, including FLAME[6], FaceScape[41], the geometry model of i3DMM[33], NPHM[66], and the previous version, ImFace[87], which represent the current state-of-the-art. The official code of FLAME is used to fit the full face scans in the test set, utilizing 300 identity parameters and 100 expression parameters. For FaceScape, we use their released bilinear model built from 938 individuals for testing. The model parameters consist of 300 identity parameters and 52 expression parameters. It is important to note that our test scans are part of the training set for FaceScape. Furthermore, we modify their official code to fit full scans, rather than just landmarks, for improved results. For i3DMM and ImFace, we retrain these models on the same training set as used for ImFace++ to ensure a fair comparison. In both cases, the identity and expression embeddings are set to 160 dimensions, which aligns with our settings. For NPHM, we utilize the model trained with the official code. To maintain consistency with our study where we focus only on the facial region, 22 facial anchors are utilized and the parameter $K_{symm}$ in NPHM is set to 8.
[…]
i3DMM[33] 320 0.759 83.87 0.938
FLAME[6] 400 0.662 88.18 0.875
FaceScape[41] 352 0.731 79.38 0.908
NPHM[66] 1000 0.637 91.67 0.934
ImFace[87] 320 0.553 94.67 0.950
ImFace++ 480 0.511 96.11 0.956
[…]
Figure 6: Reconstruction comparison with i3DMM[33], FLAME[6], FaceScape[41], NPHM[66], and ImFace[87]. Each column corresponds to a person with a non-neutral expression. Visually inspected, ImFace++ captures much richer shape variations and high-frequency details with more compact latent embeddings. Zoom in for a better view.
Figure 7: Comparison of correspondences with i3DMM[33], ImFace[87], and NPHM [66]. The leftmost face is morphed into multiple expressions (middle columns) and different identities (rightmost columns). ImFace++ surpasses the other backward-deformation methods, i.e. i3DMM and ImFace, and achieves results comparable to the forward-deformation method NPHM. Due to the forward-deformation mechanism, there are no cross-identity correspondence results for NPHM. Zoom in for a better view.
Qualitative Evaluation. Figure 6 provides a visual demonstration of the reconstruction results obtained by various models, with each row corresponding to a test individual displaying a non-neutral expression. i3DMM, being the first deep implicit model for human heads, exhibits limitations in capturing complex deformations and high-frequency details when faced with relatively intricate scenarios involving unseen facial shapes, resulting in artifacts on the reconstructed faces. FLAME and FaceScape, while effective in certain aspects, are constrained by fixed resolution and uniform mesh topology. FLAME excels at representing identity characteristics but struggles with non-linear deformations, leading to rigid facial expressions. FaceScape delivers more favourably results, primarily owing to high-quality training scans and the inclusion of test faces in the training set. However, it still struggles to accurately portray expression morphs and high-frequency details. NPHM utilizes local fields to model the SDF in a neutral expression, but the globally conditioned deformation field makes it challenging to represent fine-grained deformations and high-frequency details. Comparatively, ImFace excels in reconstructing facial shape by using the disentangled deformation fields, preserving subtle and rich nonlinear facial muscle deformations with fewer identity and expression latent parameters, such as frowns and pouts. However, it falls short of representing high-frequency details. In contrast, ImFace++ significantly improves upon its predecessor. The introduction of RDF within the template space enables ImFace++ to effectively handle intricate features, as demonstrated by its success in representing details such as eyelids (visible in the first two rows in Figure 6) and wrinkles (visible in the third to fifth rows).
Quantitative Evaluation. To ensure a fair comparison, all faces in this evaluation undergo preprocessing to remove inner structures like eyeballs. This step allows for the computation of quantitative metrics within the same facial region for all models. Specifically, the symmetric Chamfer distance, F-score, and normal consistency (N.C.) are employed as metrics, with a strict F-score threshold set at 1mm. The results are presented in Table I, clearly demonstrating that ImFace++ outperforms its counterparts by a significant margin across all metrics, unequivocally validating its effectiveness. Additionally, a color-coded distance (fit-to-scan) is used to visualize the reconstruction errors in Figure 6. Overall, ImFace++ achieves more accurate face reconstruction results compared to other methods, particularly excelling in capturing nonlinear deformations and high-frequency facial details.
Metrics EDE ($mm$${}^{{\dagger}}$ TDE ($mm$${}^{{\dagger}}$
i3DMM[33] - 4.474
NPHM[66] 3.130 -
ImFace[87] 2.805 3.116
ImFace++ 2.383 1.620
TABLE II: Quantitative correspendences comparison with the state-of-the-art methods. (${}^{{\dagger}}$ower is better.)
4.3.2 Correspondence
In contrast to existing methods that typically rely on accurate face registration, ImFace++ leverages the disentangled deformation fields to automatically learn correspondences among different face scans. Furthermore, we incorporate the training critic, specifically the landmark consistency loss ${\cal L}_{lmk_{c}}^{i}$, to enhance this feature. This evaluation is designed to assess the correspondence establishment capabilities of various INR-based models. In particular, given two 3D faces, we use i3DMM, ImFace, and ImFace++ to fit them, subsequently deforming the densely sampled points to the template space. This process ensures point-to-point correspondences, which are achieved through a nearest neighbor search using the KD tree algorithm for efficiency. Considering that NPHM is a forward-deformation-based method, we deform the sampled points in a neutral expression to another expression and examine the correspondences.
Qualitative Evaluation. Figure 7 provides a visual comparison of the results, with manually painted color patterns on the shapes to facilitate quality assessment. In alignment with our approach, i3DMM incorporates a landmark consistency loss to enhance correspondences. However, it employs a single deformation field to represent both expression and identity deformations in an intertwined manner. Consequently, the correspondences may not be accurately learned in certain areas, such as the mouth (regions F4-F5 in the painted color patterns), forehead (regions B3-B6), and nostril (regions E4-E5). ImFace successfully establishes correspondences across various expressions and identities. Nonetheless, it is noticeable that minor internal texture discrepancies occasionally occur around the mouth (regions F4-F5, G4-G5). This is primarily because facial shapes undergo significant changes in these localized areas with different expressions. ImFace++ addresses this challenge by employing densely sampled landmarks (increasing from 68 in ImFace to 4,096 in ImFace++) to supervise landmark consistency loss during training, resulting in more accurate outcomes. Due to the forward-deformation pipeline, NPHM naturally exhibits more consistent correspondences than backward-deformation approaches. However, it may overlook exaggerated expressions like an open mouth, potentially leading to semantically incorrect correspondences (regions G4-G5). In contrast, ImFace++ achieves comparable or even superior results to NPHM, demonstrating its effectiveness.
Figure 8: Module analysis. The disentangled deformation fields are responsible for learning deformations related to facial expressions and identities. The refinement displacement field enables the representation of high-frequency details.
Quantitative Evaluation. To further validate the effectiveness of our method, we conduct a quantitative evaluation on the learned correspondences similar to i3DMM [33]. For each test face $F$ with identity $i$ from FaceScape, we first identify the face $F_{ne}$ with the same identity but in a neutral expression. We then upsample the registered mesh resolution and sample 22,481 points from faces $F$, $F_{ne}$, and the template face, which are not used during training, serving as the ground-truth annotations for correspondences. We use i3DMM, ImFace, and ImFace++ to deform the points sampled from face $F$ to the canonical space and the template space, respectively. We calculate the average Expression Deformation Error (EDE), which quantifies the errors between the deformed points and the ground-truth points in the canonical space, and the Total Deformation Error (TDE), which measures the errors in the template space. Note that i3DMM employs a single deformation field to represent both expression and identity deformations, allowing it to compute only the TDE metric. On the other hand, NPHM models only the deformation of expression and we can calculate only the EDE metric. For NPHM, we follow the approach of SNARF [31] and utilize iterative root finding to locate the canonical positions of the points sampled from face $F$, allowing us to compute the EDE metric. As shown in Table II, ImFace++ outperforms the other methods in terms of correspondence quality. Notably, it even surpasses the forward-deformation method, NPHM, underscoring the effectiveness of the disentangled deformation fields, which play a crucial role in addressing the challenges associated with learning the RDF. It is worth mentioning that the results in Table II also encompass the reconstruction errors.
4.4 Representation Ability
To further assess the representation ability of ImFace++, we conduct comprehensive evaluations in both the observed and latent representation spaces. These evaluations comprise three main aspects. i) Module Analysis: we analyze the learned content of the various modules introduced in ImFace++. ii) Compact Embedding Space Interpolation: given the compact embedding space of ImFace++, we perform animations on the generated or reconstructed facial shapes through interpolation in the identity, expression, and detail embeddings. iii) Variation Visualization: we apply Principal Components Analysis (PCA) to the learned expression and identity embeddings to visualize model variations. Additionally, we employ t-SNE [85] to assess the unsupervised learning ability of ImFace++.
Module Analysis. Figure 8 presents the reconstructed test scans by ImFace++, offering insights into the content learned by the different modules proposed in our method. In the third and fourth columns, we can observe that the disentangled deformation fields, managed by ExpNet for expression and IDNet for identity, successfully capture rich and complex deformations. However, the reconstructed facial shape in these cases lacks high-frequency details. Nevertheless, after training the RDF within the template space, ImFace++ is capable of representing facial shapes that incorporate personalized high-frequency details. This is particularly evident in features like the eyelids in the first two rows and the mouth shape in the third row. In the ablation study section, we offer further evidence and a more comprehensive analysis.
Compact Embedding Space Interpolation. The seamless animation of facial shapes through interpolation in the identity, expression, and detail embeddings demonstrates that ImFace++ has learned compact latent embedding spaces. As illustrated in Figure 9, our identity embeddings encompass the overall shape, including the positioning of the nose, eyes, and mouth, as well as the gender of the face. In contrast, the expression embeddings capture deformations specific to facial components. Furthermore, when comparing the first and last columns in Figure 9, it becomes evident that the detail embeddings represent personalized high-frequency details of the face, such as wrinkles and dimples. As demonstrated, even when the embeddings used for interpolation exhibit significant changes over a wider range (in the last two rows), our method is capable of generating smooth and coherent results.
Figure 9: Embedding interpolation results. The Trainset and Testset scans can be smoothly animated by interpolating in the embedding space.
Figure 10: Visualization of the expression and identity principal components of ImFace++ on FaceScape [41].
Figure 11: Visualizing the distributions of high-dimensional expression and identity embeddings using t-SNE. (a) The embedding distribution of 6 typical expressions from the training set. (b) The identity embedding distribution from the test set, which involves 16 individuals.
Variation Visualization. We apply PCA to the learned expression and identity embeddings to visualize the model capacity, as depicted in Figure 10. The standard deviations for the expression and identity embeddings are set to ±$\pm~{}0.25$ and ±$\pm~{}5.0$, respectively. The embeddings encoding facial details are fixed at the mean value calculated from the entire training set. In particular, four expression principal components are visualized in Figure 10 (a). Despite significant expression changes, the faces maintain a consistent identity. ImFace++ is capable of producing vivid expressions by learning expression components from thousands of unique embeddings. In Figure 10 (b), a similar phenomenon is observed in the learned identity components, where facial expressions remain stable while identity varies. These experimental results demonstrate that a good disentanglement between expression and identity has been achieved, which is crucial for generating novel faces by reweighting the singular values.
Additionally, we employ t-SNE[85] to visualize the learned high-dimensional expression embeddings from 4,992 training scans, encompassing six typical expressions, as depicted in Figure 11 (a). It is evident that our network can effectively distinguish different expression types through unsupervised learning, solely from expression-related shape morphs, without relying on explicit expression labels. Furthermore, Figure 11 (b) presents the visualization of identity embeddings, involving 319 face scans from 16 individuals in the test set. Upon visual inspection, our model successfully captures inter-subject differences even under various complicated expressions.
4.5 Ablation Study
ImFace++ is constructed upon the following core components: Disentangled Deformation Fields (DDF), Refinement Displacement Field (RDF), Neural Blend-Field (Blend), and Improved Expression Embedding Learning (Extend). We experimentally verify the credits of these designs. When performing ablation of RDF, the model is essentially downgraded to ImFace. During its training process, we increase the number of the sampled points to 4,096 to supervise the landmark consistency loss, ensuring a fair comparison with ImFace++.
[…]
(a) w/o DDF 0.568 94.59 0.950
(b) w/o Blend 0.564 93.41 0.949
(c) w/o Extend 0.596 91.75 0.947
(d) $*{}^{*}$ (w/o RDF) 0.547 95.33 0.952
(e) ImFace++ 0.511 96.11 0.956
TABLE III: Quantitative ablation study results. $*{}^{*}$ denotes the use of 4,096 landmarks to supervise the landmark consistency loss during model training.
On Disentangled Deformation Fields. To underscore the significance of the disentangled deformation learning process, we establish a baseline network that incorporates only one deformation field to universally learn face shape morphs. In this configuration, ${\mathbf{z}}_{exp}$ and ${\mathbf{z}}_{id}$ are concatenated and used as the input for the hyper net. Figure 12 (a) visually demonstrates the results. Despite some fine-grained details introduced by other designs, there is noticeable chaos in the reconstructed faces, especially for those with significant expressions. The quantitative results presented in Table III also indicate the significance of decoupled deformation learning.
On Neural Blend-Field. We substitute the Neural Blend-Field in $\mathcal{E}$, $\mathcal{I}$, $\mathcal{T}$ with standard MLPs containing an equivalent number of parameters. These MLPs directly predict global deformations or SDF values for the entire face. As illustrated in Figure 12 (b), visible blurriness appears due to the limited capability to learn non-linear facial changes. The quantitative results in Table III further affirm the necessity of the Neural Blend-Field in achieving accurate representations.
On Improved Embedding Learning. This strategy is introduced to facilitate the learning of more diverse and fine-grained facial expression deformations. As demonstrated in Figure 12 (c), when the number of expression embeddings is constrained to be the same as the expression categories, the generated expressions tend to be average. Furthermore, for exaggerated expressions, such as mouth stretching, the counterpart models struggle to converge to a reasonable state.
On Refinement Displacement Field. RDF in ImFace++ is designed to learn and represent high-frequency details in facial shapes. A direct comparison between Figure 12 (d) and (e) illustrates that while ImFace can partially capture nonlinear facial deformations, it faces challenges in representing high-frequency details and personalized nuances. In contrast, ImFace++ excels at learning these intricate, subtle differences, by specially encoding them within the template space, exemplified by the sophisticated reproduction of features like the eyelid and wrinkles around the mouth. Furthermore, the quantitative validation presented in Table III underscores the crucial role of this newly introduced RDF module in ImFace++ for constructing a highly convincing 3D morphable model.
Figure 12: Qualitative ablation study results.
Figure 13: Expression editing results. The faces present in the first row provide the source identity and facial detail embeddings.
4.6 Applications
4.6.1 Expression Editing
Similar to existing 3DMMs, ImFace++ serves as a comprehensive facial representation built upon prior distributions of facial expression and identity morphs, making it suitable for a variety of downstream applications. In Figure 13, we present expression editing results obtained from the FaceScape test set. The faces in the first row are the reconstructed ones providing source identities and neutral expressions, while the faces in the first column showcase the target expressions. In this figure, the first four rows depict reconstructed faces from the test set, and the last three rows feature generated faces from the train set. It is evident that facial expression editing can be easily accomplished by modifying the expression embeddings. These lifelike 3D faces demonstrate the impressive representational capabilities of ImFace++.
4.6.2 Facial Detail Transfer
In Fig. 14, ImFace++ demonstrates its capability to transfer high-frequency facial details across different identities and expressions. In particular, the target faces in columns 4 to 7 exhibit various eyelid types, a challenging task for explicit face representations. ImFace++ accomplishes the transfer of high-frequency details with impressive fidelity by merely swapping the detail embeddings. Furthermore, the model can also transfer facial wrinkles, such as frown lines and crow’s feet in the last two columns, to the source faces, showcasing its ability to capture and replicate fine-grained facial features with remarkable precision.
5 Conclusion
This paper introduces ImFace++, a sophisticated nonlinear 3D morphable face model with INRs. ImFace++ excels in capturing complex facial shape variations through the use of two explicitly disentangled deformation fields associated with expression- and identity-variation, along with a refinement displacement field that encodes high-frequency facial geometry details, contributing to the creation of highly realistic facial representations. To enhance the precision in capturing detailed facial deformations and geometries, ImFace++ also introduces a Neural Blend-Field. Furthermore, an improved embedding learning strategy is proposed that enables more fine-grained expressions. Additionally, the paper presents an effective preprocessing pipeline that enables INRs to work with non-watertight facial surfaces for the first time. Experimental results demonstrate that ImFace++ is highly competent and outperforms the state-of-the-art methods in this domain.
Figure 14: Facial detail transfer results. ImFace++ is able to edit high-frequency facial details by modifying the corresponding embeddings. The faces in the first column provide the source identities and expression embeddings.
While this work represents a significant advancement in 3D facial shape representation, the focus has primarily been on face geometry modeling, with less consideration given to facial appearance. Although a basic appearance model can be achieved by adding a color field, there leaves space for further exploration in developing a more comprehensive INR-based 3DMM that accurately represents facial appearance with realistic diffuse and specular albedo. Addressing this issue could be a significant direction for future research.
Acknowledgments
This work is partly supported by the National Key R&D Program of China (2022ZD0161902), the National Natural Science Foundation of China (No. 62176012, 62202031, 62022011), Beijing Municipal Natural Science Foundation (No. 4222049), and the Fundamental Research Funds for the Central Universities.